OpenStudy (anonymous):

Complete the standard form of the equation: y = 10- 5/2x ____x + ____y = 20

OpenStudy (anonymous):

Please, can someone help? :(

OpenStudy (anonymous):

:) So, since this equation (____x + ____y = 20) has x & y on the same side, the first thing you should do is bring the x-term from to the other side: so you have y + (5/2)x = 10 then multiply both sides by 2 to get 20 on the right side 2( y + (5/2)x) = 10(2) & you get 2y + 5x = 20

OpenStudy (anonymous):

Thank you so much! ^-^